# TRABSA: Interpretable Sentiment Analysis of Tweets using Attention-based BiLSTM and Twitter-RoBERTa

### v1

@@ 4.1 Experimental Setup
4.2 Evaluation Metrics
We use a variety of assessment indicators in our research to determine our models’ overall success. These measurements shed light on how well the models categorize sentiments correctly across various classifications. The F1-score, accuracy, precision, recall, and macro-average metrics are among the crucial assessment measures that are employed. These are computed using the following definitions of true positives (TP), false positives (FP), and false negatives (FN):
@@ 4.2 Evaluation Metrics
Together, these evaluation measures offer a thorough analysis of the performance of the model employed in this research, assisting in the choice and improvement of SA models.
4.3 Results
Analyzing the performance of various models across different word embedding techniques reveals significant variations in their effectiveness for SA tasks. Among the traditional models, GBM and LightGBM consistently demonstrate strong performance across all word embedding techniques, achieving macro average F1-scores ranging from 79% to 83%. Pre-trained transformer models like BERT and RoBERTa exhibit competitive performance, with the RoBERTa model showcasing high accuracy across various configurations, especially when combined with advanced neural network architectures. Notably, the TRABSA model outperforms all other models across all evaluation metrics, demonstrating exceptional macro average precision (94%), macro average recall (93%), macro average F1-score (94%), and accuracy of 94%. This significant improvement underscores the efficacy of the TRABSA framework in SA tasks, surpassing even state-of-the-art transformer models like RoBERTa.
@@ 4.3 Results
Compared to the best-performing traditional models like LightGBM (83%) and advanced transformer models like RoBERTa (84%), the TRABSA model achieves an impressive accuracy of 94%, indicating a substantial improvement of at least 11% over the closest competitors. The TRABSA model consistently outperforms others in accuracy and macro average precision, recall, and F1-score, demonstrating its robustness and effectiveness across different evaluation criteria. Even when compared to sophisticated neural network architectures like Single Hidden Layer NN (72%) and 3 Hidden Layers NN (70%), the TRABSA model exhibits a remarkable performance boost of over 20%, further emphasizing its superiority in SA tasks. The TRABSA model’s outstanding performance underscores its hybrid architecture’s efficacy, which integrates transformer-based mechanisms, attention mechanisms, and BiLSTM networks to capture nuanced sentiment patterns effectively. The findings suggest that while pre-trained transformer models like BERT and RoBERTa offer competitive performance, customized architectures like TRABSA tailored specifically for SA tasks can yield substantial accuracy and predictive power improvements.
4.4 Robustness Test of TRABSA Model
@@
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = r=2, cell13 = r=2, cell14 = c=4, cell31 = r=5, cell32 = r=5, cell81 = r=5, cell82 = r=5, cell131 = r=5, cell132 = r=5, cell181 = r=5, cell182 = r=5, cell231 = r=5, cell232 = r=5, cell281 = r=5, cell282 = r=5, vlines, hline1,3,8,13,18,23,28,33 = -, hline2 = 4-7, hline4-7,9-12,14-17,19-22,24-27,29-32 = 3-7, Dataset Type & Dataset Name DL Models Evaluation Metrics
@@ 3 Hidden Layers NN 85% 83% 84% 85%
BiLSTM+3 Hidden Layers NN 87% 86% 86% 87%
@@
Proposed TRABSA Model 85% 85% 85% 86%
@@ 3 Hidden Layers NN 92% 92% 92% 92%
External Reddit Dataset Single Hidden Layer NN 94% 94% 94% 95%
@@ 3 Hidden Layers NN 94% 94% 94% 95%
Proposed TRABSA Model 94% 94% 94% 95%
The TRABSA model consistently demonstrates robustness and generalizability across datasets and DL architectures, as evidenced by its consistent performance metrics. Across various datasets, including the Global COVID-19 Dataset, the USA COVID-19 Dataset, the External Twitter Dataset, the Reddit Dataset, the Apple Dataset, and the US Airline Dataset, the TRABSA model consistently achieves high macro average precision, recall, F1-score, and accuracy values. For instance, in the Global COVID-19 Dataset, the TRABSA model attains an impressive 98% macro average precision, recall, F1-score, and accuracy. Similarly, the USA COVID-19 dataset maintains high scores, averaging 85% across all metrics. The trend continues across External datasets, with the TRABSA model consistently performing exceptionally well, achieving an average accuracy of 97% on the Twitter Dataset, 94% on the Reddit Dataset, 90% on the Apple Dataset, and 95% on the US Airline Dataset. These consistent and high-performance metrics underscore the reliability and effectiveness of the TRABSA model across diverse datasets and DL architectures, reaffirming its robustness and generalizability in SA tasks.
5 Interpretability Analysis
@@
5.1 SHAP
A useful tool for deciphering and understanding the results of ML models is the SHapley Additive exPlanations (SHAP) framework [Jahin et al., 2023]. The computation and presentation of the relevance assigned to each characteristic in the prediction process are made easier by utilizing the SHAP Python package. The calculation of SHAP values, which act as measurements of feature contribution and improve the interpretability of ML models, is essential to the SHAP framework. When the features ($x$) are unknown, a SHAP value specifies how to go from the expected or base value $E[f(x)]$ to the actual output $f$. Furthermore, by clarifying the direction of the link between features and the target variable, SHAP values shed light on how characteristics affect predictions. A characteristic with a SHAP value of 1 or -1, for example, significantly influences the prediction for a given data point in a favorable or negative way. On the other hand, a feature that approaches 0 in SHAP value has a negligible contribution to the prediction [Jahin et al., 2023]. A range of graphs are provided by the SHAP framework to help in the understanding of feature contributions and to aid in the interpretation and justification of ML models. The following represents how SHAP values are calculated:
@@ 5.1 SHAP
5.1.1 SHAP Text Plot
(a) Positive sentiments
(b) Neutral sentiments
(c) Negative sentiments
@@ 5.1.1 SHAP Text Plot
A thorough illustration of how specific tokens inside a text instance affect a TRABSA model’s output may be seen in Figure 9. The graphic illustrates sections that either boost (in red) or reduce (in blue) the model’s sentiment forecast by superimposing significance values over the original text. This makes it possible to comprehend how certain words or phrases fit into the larger feeling that the text is trying to convey in a nuanced way. Furthermore, the hierarchical structure of significance values preserves the structural linkages between tokens, providing insights into intricate interactions within the text. A comprehensive picture of how the text’s combined characteristics affect the model’s output is given by the force plot that goes with it; positive features raise the prediction while negative features reduce it. Interactive capability further improves interpretability by allowing users to investigate the connections between text segments and how those connections affect the model’s predictions. All things considered, the figure makes it easier to comprehend how the TRABSA model makes decisions and helps with text-based data interpretation and analysis.
5.1.2 SHAP Bar Plot
@@
The NLP global summaries of the influence of tokens inside a dataset are shown in Figure 10. Every bar in the graphic illustrates how significant a token is overall in affecting the model’s predictions over the full dataset. Each bar’s height represents the influence of the token; higher bars denote greater significance. The plot collects the individual contributions of tokens over numerous instances by compressing the Explanation object over its rows, usually by summing. This method generates a bar chart with as many columns as there were unique tokens in the original dataset by treating each kind of token as a feature. Big groups are split up, and each token gets an equal portion of the total group significance value if the Explanation object contains hierarchical values. Furthermore, arranging the bar chart in a descending sequence helps reveal which tokens have a major impact on the model’s predictions.
5.2 LIME
A methodical technique for evaluating individual predictions given by complicated ML models is provided by the LIME method, which stands for Locally Interpretable Model-Agnostic Explanations [Jahin et al., 2023]. This approach works by estimating the model’s behavior around a given forecast. Fundamentally, LIME utilizes a local linear explanation model that follows Equation 9 to the letter, making it an additive feature attribution method. LIME presents the idea of ”interpretable inputs,” which are condensed representations of the original inputs and are represented as $x_{0}$. A binary vector of interpretable inputs is mapped to the original input space via the transformation $x=h_{x}(x_{0})$. Different kinds of mappings $h_{x}$ are used for different input spaces. For example, when applied to bag-of-words text characteristics, $h_{x}$ translates a binary vector (signaling the presence or absence of words) to the appropriate word count in the source text. LIME aims to minimize the objective function to determine the coefficients $\phi$:
@@ 5.2 LIME
5.2.1 LIME Text Explainer Plot
@@
A useful tool for understanding how particular elements or tokens inside a text affect the model’s predictions is the LIME Text Explainer visualization. Every text token is plotted along the horizontal axis in Figure 11, and its contribution to the prediction is indicated along the vertical axis. Usually, the plot shows how important tokens are by emphasizing how they affect the model’s output. Analyzing each token’s contribution amount and direction is necessary to understand a LIME Text Explainer plot. Stronger impacts on the model’s prediction are indicated by tokens with bigger positive or negative contributions. On the other hand, tokens with contributions closer to zero indicate a negligible influence on the prediction. Additionally, the plot could draw attention to particular words or phrases that greatly impact the model’s ability to make decisions. Understanding these influential tokens can provide valuable insights into how the model processes and evaluates textual data.
6 Discussions
@@
Our methodology and results demonstrate a significant advancement in SA compared to existing literature. While prior studies have explored diverse models and techniques, our TRABSA model introduces a unique hybrid approach combining transformer-based architectures, attention mechanisms, and BiLSTM networks. This innovative combination enables our model to effectively capture nuanced sentiment patterns, resulting in notably higher accuracy and performance across multiple evaluation metrics than traditional and state-of-the-art transformer models such as BERT and RoBERTa. We conducted a thorough analysis to assess the robustness of our TRABSA model across various datasets and scenarios, consistently observing superior performance across multiple datasets, including extended and external ones. Additionally, our model demonstrates resilience to variations in sentiment expression and context, reaffirming its reliability in diverse real-world scenarios.
@@
This novel hybrid approach offers several benefits to the field, including unparalleled accuracy, robustness, and generalizability across diverse datasets and scenarios. By leveraging the strengths of each component, the TRABSA model can revolutionize SA applications, providing researchers, businesses, and policymakers with deeper insights into public opinion, consumer sentiment, and social trends. Its innovative architecture and superior performance represent a significant advancement in the quest for more accurate and reliable SA tools, with implications extending beyond academic research.
@@
The practical implications of the TRABSA model’s advancements are profound, offering tangible benefits across various real-world applications. In market research, the model’s ability to accurately analyze sentiment from social media, customer reviews, and other online sources empowers companies to gain valuable insights into consumer preferences, market trends, and brand sentiment. This knowledge informs strategic decision-making processes, product development strategies, and marketing campaigns, ultimately enhancing customer satisfaction and competitive advantage. Furthermore, in social media monitoring and reputation management, the TRABSA model equips organizations with tools to monitor public sentiment, identify emerging issues or crises, and proactively respond to real-time customer feedback. Detecting and addressing potential issues early on enables businesses to safeguard their reputation and maintain positive relationships with their target audience. Additionally, in the context of public opinion analysis and political discourse, the TRABSA model provides policymakers and analysts with a powerful tool for gauging public sentiment, identifying key concerns, and tracking changes in public perception over time. This knowledge informs policy decisions, communication strategies, and crisis management efforts, ultimately contributing to more informed and responsive governance. The practical applications of the TRABSA model extend across a wide range of industries and domains, offering transformative benefits for businesses, governments, and society as a whole.
7 Conclusions and Future Directions
@@
Our research has yielded significant findings and contributions to SA. We have achieved remarkable results by developing and evaluating the TRABSA model, a novel hybrid approach combining transformer-based architectures, attention mechanisms, and BiLSTM networks. Leveraging the latest RoBERTa-based transformer model and expanding the datasets, we have demonstrated the TRABSA model’s exceptional accuracy and relevance, bridging existing gaps in SA benchmarks. Thorough comparisons of word embedding techniques and methodical labeling of tweets using lexicon-based approaches have further enhanced the effectiveness of SA methodologies. Our experiments and benchmarking efforts have highlighted the superiority of the TRABSA model over traditional and state-of-the-art models, showcasing its versatility and robustness across diverse datasets and scenarios. With macro-average precision of 94%, macro-average recall of 93%, macro-average F1-score of 94%, and accuracy of 94%, our model has proven its efficacy in capturing nuanced sentiment patterns. Additionally, exploring model interpretability techniques using SHAP and LIME has enhanced our understanding and trust in the TRABSA model’s predictions, reinforcing its practical applicability.
@@
Despite the significant advancements achieved in our research, several avenues remain for future exploration and improvement in interpretable SA. Firstly, there is scope for refining and expanding model interpretability techniques to provide deeper insights into the factors influencing sentiment predictions. Additionally, integrating multimodal data sources such as text, images, and audio could enhance the richness and accuracy of SA. Addressing ethical considerations regarding bias, fairness, and privacy in SA models is paramount for responsible deployment and usage. Furthermore, exploring the application of SA in emerging domains such as healthcare, finance, and politics could uncover new challenges and opportunities for research and innovation. Overall, continued research in interpretable SA holds the potential to drive meaningful advancements in AI technologies and contribute to more informed decision-making in various fields.
8 Data Availability
@@
The extended datasets, comprising the Global Twitter COVID-19 Dataset and the USA Twitter COVID-19 Dataset, are publicly available for download from the Extended Covid Twitter Datasets repository [Jahin, 2023]. Additionally, the external datasets used in our research were sourced from Kaggle, including the Twitter Dataset, Reddit Dataset, Apple Twitter Dataset, and US Airline Twitter Dataset.

### v2

@@ Results
Evaluation Metrics
We use a variety of assessment indicators in our research to determine our models’ overall success. These measurements show how well the models correctly categorize sentiments across various classifications. The F1-score, accuracy, precision, recall, and macro-average metrics are among the crucial assessment measures employed. These are computed using the following definitions of true positives (TP), false positives (FP), and false negatives (FN):
@@
Together, these evaluation measures offer a thorough analysis of the model’s performance employed in this research, assisting in the choice and improvement of SA models.
Results
Analyzing the performance of various models across different word embedding techniques reveals significant variations in their effectiveness for sentiment analysis tasks. Among the traditional models, GBM and LightGBM consistently demonstrate strong performance across all word embedding techniques, achieving macro average F1-scores ranging from 79% to 83%. Pre-trained transformer models like BERT and RoBERTa exhibit competitive performance, with the RoBERTa model showcasing high accuracy across various configurations, especially when combined with advanced neural network architectures. Notably, the TRABSA model outperforms all other models across all evaluation metrics, demonstrating exceptional macro average precision (94%), macro average recall (93%), macro average F1-score (94%), and accuracy of 94% (see Table LABEL:tab3). This significant improvement underscores the efficacy of the TRABSA framework in sentiment analysis tasks, surpassing even state-of-the-art transformer models like RoBERTa.
We used BoW, TFIDF, word2vec, BERT, SBERT, and RoBERTa as word embeddings with all the ML and two DL models. When we noticed a significant improvement in the accuracy (above 80%) while using RoBERTa word embeddings with simple DL architectures, we decided to implement more complex hybrid DL models: BiLSTM+3 Hidden Layers NN, BiLSTM+CNN, and TRABSA.
@@
Compared to the best-performing traditional models like LightGBM (83%), stacked LGBM+KNN+MLP, and advanced hybrid state-of-the-art BiLSTM+3 Hidden Layers NN (85%), the TRABSA model achieves an impressive accuracy of 94%, indicating a substantial improvement of at least 11% over the closest competitors. The TRABSA model consistently outperforms others in accuracy and macro average precision, recall, and F1-score, demonstrating its robustness and effectiveness across different evaluation criteria. Even when compared to sophisticated neural network architectures like Single Hidden Layer NN (84%) and 3 Hidden Layers NN (84%), the TRABSA model exhibits a remarkable performance boost of 11.90%, further emphasizing its superiority in sentiment analysis tasks. The TRABSA model’s outstanding performance underscores its hybrid architecture’s efficacy, which integrates transformer-based mechanisms, attention mechanisms, and BiLSTM networks to capture nuanced sentiment patterns effectively. The findings suggest that while pre-trained transformer models like BERT and RoBERTa offer competitive performance, customized architectures like TRABSA tailored specifically for sentiment analysis tasks can yield substantial accuracy and predictive power improvements.
Robustness Test of TRABSA Model
@@
Dataset Type Dataset Name DL Models Evaluation Metrics
@@ 3 Hidden Layers NN 85% 83% 84% 85%
BiLSTM+3 Hidden Layers NN 85% 85% 85% 86%
@@
Proposed TRABSA Model 87% 86% 86% 87%
@@ 3 Hidden Layers NN 92% 92% 92% 92%
External Reddit Dataset Single Hidden Layer NN 94% 93% 94% 94%
@@ 3 Hidden Layers NN 94% 93% 93% 94%
Proposed TRABSA Model 95% 95% 95% 96%
The TRABSA model consistently demonstrates robustness and generalizability across datasets and DL architectures, as evidenced by its consistent performance metrics. Across various datasets, including the Global COVID-19 Dataset, the USA COVID-19 Dataset, the External Twitter Dataset, the Reddit Dataset, the Apple Dataset, and the US Airline Dataset, the TRABSA model consistently achieves high macro average precision, recall, F1-score, and accuracy values. For instance, in the Global COVID-19 Dataset, the TRABSA model attains an impressive 98% macro average precision, recall, F1-score, and accuracy. Similarly, the USA COVID-19 dataset maintains high scores, obtaining 87% in terms of accuracy. The trend continues across External datasets, with the TRABSA model consistently performing exceptionally well, achieving an average accuracy of 97% on the Twitter Dataset, 95% on the Reddit Dataset, 90% on the Apple Dataset, and 96% on the US Airline Dataset (see Table 4). These consistent and high-performance metrics underscore the reliability and effectiveness of the TRABSA model across diverse datasets and DL architectures, reaffirming its robustness and generalizability in SA tasks.
Interpretability Analysis
@@
SHAP
A useful tool for deciphering and understanding the results of ML models is the SHapley Additive exPlanations (SHAP) framework [43]. The computation and presentation of the relevance assigned to each characteristic in the prediction process are made easier by utilizing the SHAP Python package. Calculating SHAP values, which measure feature contribution and improve the interpretability of ML models, is essential to the SHAP framework. When the features ($x$) are unknown, a SHAP value specifies how to go from the expected or base value $E[f(x)]$ to the actual output $f$. Furthermore, by clarifying the direction of the link between features and the target variable, SHAP values shed light on how characteristics affect predictions. A characteristic with a SHAP value of 1 or -1, for example, significantly influences the prediction for a given data point favorably or negatively. On the other hand, a feature that approaches 0 in SHAP value has a negligible contribution to the prediction [43]. A range of graphs are provided by the SHAP framework to help in the understanding of feature contributions and to aid in the interpretation and justification of ML models. The following represents how SHAP values are calculated:
@@
SHAP Text Plot
(a) Positive sentiment
(b) Neutral sentiment
(c) Negative sentiment
@@
A thorough illustration of how specific tokens inside a text instance affect a TRABSA model’s output may be seen in Figure 8. The graphic illustrates sections that boost (in red) or reduce (in blue) the model’s sentiment forecast by superimposing significance values over the original text. This makes it possible to comprehend how certain words or phrases fit into the larger feeling the text is trying to convey nuancedly. Furthermore, the hierarchical structure of significance values preserves the structural linkages between tokens, providing insights into intricate interactions within the text. A comprehensive picture of how the text’s combined characteristics affect the model’s output is given by the force plot that goes with it; positive features raise the prediction while negative features reduce it. By allowing users to investigate the connections between text segments and how those connections affect the model’s predictions, interactive capability further improves interpretability. All things considered, the figure makes it easier to comprehend how the TRABSA model makes decisions and helps with text-based data interpretation and analysis.
SHAP Bar Plot
@@
The NLP global summaries of the influence of tokens inside a dataset are shown in Figure 9. Every bar in the graphic illustrates how significantly a token affects the model’s predictions over the full dataset. Each bar’s height represents the influence of the token; higher bars denote greater significance. The plot collects the individual contributions of tokens over numerous instances by compressing the Explanation object over its rows, usually by summing. This method generates a bar chart with as many columns as there were unique tokens in the original dataset by treating each kind of token as a feature. Big groups are split up, and each token gets an equal portion of the total group significance value if the Explanation object contains hierarchical values. Furthermore, arranging the bar chart in a descending sequence helps reveal which tokens have a major impact on the model’s predictions.
LIME
The LIME method provides a methodical technique for evaluating individual predictions given by complicated ML models, which stands for Locally Interpretable Model-Agnostic Explanations [43]. This approach works by estimating the model’s behavior around a given forecast. Fundamentally, LIME utilizes a local linear explanation model that follows Equation 9 to the letter, making it an additive feature attribution method. LIME presents the idea of “interpretable inputs," which are condensed representations of the original inputs and are represented as $x_{0}$. A binary vector of interpretable inputs is mapped to the original input space via the transformation $x=h_{x}(x_{0})$. Different kinds of mappings $h_{x}$ are used for different input spaces. For example, when applied to bag-of-words text characteristics, $h_{x}$ translates a binary vector (signaling the presence or absence of words) to the appropriate word count in the source text. LIME aims to minimize the objective function to determine the coefficients $\phi$:
@@
LIME Text Explainer Plot
@@
The LIME Text Explainer visualization is a useful tool for understanding how particular elements or tokens inside a text affect the model’s predictions. Every text token is plotted along the horizontal axis in Figure 10, and its contribution to the prediction is indicated along the vertical axis. Usually, the plot shows how important tokens are by emphasizing how they affect the model’s output. Analyzing each token’s contribution amount and direction is necessary to understand a LIME Text Explainer plot. Stronger impacts on the model’s prediction are indicated by tokens with bigger positive or negative contributions. On the other hand, tokens with contributions closer to zero indicate a negligible influence on the prediction. Additionally, the plot could draw attention to particular words or phrases that greatly impact the model’s ability to make decisions. Understanding these influential tokens can provide valuable insights into how the model processes and evaluates textual data.
Discussions
@@
Our methodology and results demonstrate a significant advancement in sentiment analysis compared to existing literature. While prior studies have explored diverse models and techniques, our TRABSA model introduces a unique hybrid approach combining transformer-based architectures, attention mechanisms, and BiLSTM networks. This innovative combination enables our model to effectively capture nuanced sentiment patterns, resulting in notably higher accuracy and performance across multiple evaluation metrics than traditional and state-of-the-art transformer models such as BERT and RoBERTa. We conducted a thorough analysis to assess the robustness of our TRABSA model across various datasets and scenarios, consistently observing superior performance across multiple datasets, including extended and external ones. Additionally, our model demonstrates resilience to variations in sentiment expression and context, reaffirming its reliability in diverse real-world scenarios.
@@
This novel hybrid approach offers several benefits to the field, including unparalleled accuracy, robustness, and generalizability across diverse datasets and scenarios. By leveraging the strengths of each component, the TRABSA model can revolutionize sentiment analysis applications, providing researchers, businesses, and policymakers with deeper insights into public opinion, consumer sentiment, and social trends. Its innovative architecture and superior performance represent a significant advancement in the quest for more accurate and reliable sentiment analysis tools, with implications extending beyond academic research.
@@
The practical implications of the TRABSA model’s advancements are profound, offering tangible benefits across various real-world applications. In market research, the model’s ability to accurately analyze sentiment from social media, customer reviews, and other online sources empowers companies to gain valuable insights into consumer preferences, market trends, and brand sentiment. This knowledge informs strategic decision-making processes, product development strategies, and marketing campaigns, ultimately enhancing customer satisfaction and competitive advantage. Furthermore, in social media monitoring and reputation management, the TRABSA model equips organizations with tools to monitor public sentiment, identify emerging issues or crises, and proactively respond to customer feedback in real-time. Detecting and addressing potential issues early on enables businesses to safeguard their reputation and maintain positive relationships with their target audience. Additionally, in the context of public opinion analysis and political discourse, the TRABSA model provides policymakers and analysts with a powerful tool for gauging public sentiment, identifying key concerns, and tracking changes in public perception over time. This knowledge informs policy decisions, communication strategies, and crisis management efforts, ultimately contributing to more informed and responsive governance. The practical applications of the TRABSA model extend across a wide range of industries and domains, offering transformative benefits for businesses, governments, and society as a whole.
Conclusions and Future Directions
@@
Our research has yielded significant findings and contributions to sentiment analysis. We have achieved remarkable results by developing and evaluating the TRABSA model, a novel hybrid approach combining transformer-based architectures, attention mechanisms, and BiLSTM networks. Leveraging the latest RoBERTa-based transformer model and expanding the datasets, we have demonstrated the TRABSA model’s exceptional accuracy and relevance, bridging existing gaps in sentiment analysis benchmarks. Thorough comparisons of word embedding techniques and methodical labeling of tweets using lexicon-based approaches have further enhanced the effectiveness of sentiment analysis methodologies. Our experiments and benchmarking efforts have highlighted the superiority of the TRABSA model over traditional and state-of-the-art models, showcasing its versatility and robustness across diverse datasets and scenarios. With macro-average precision of 94%, macro-average recall of 93%, macro-average F1-score of 94%, and accuracy of 94%, our model has proven its efficacy in capturing nuanced sentiment patterns. Additionally, exploring model interpretability techniques using SHAP and LIME has enhanced our understanding and trust in the TRABSA model’s predictions, reinforcing its practical applicability.
@@
Despite the significant advancements achieved in our research, several avenues remain for future exploration and improvement in interpretable sentiment analysis. Firstly, there is scope for refining and expanding model interpretability techniques to provide deeper insights into the factors influencing sentiment predictions. Additionally, integrating multimodal data sources such as text, images, and audio could enhance the richness and accuracy of sentiment analysis. Addressing ethical considerations regarding bias, fairness, and privacy in sentiment analysis models is paramount for responsible deployment and usage. Furthermore, exploring the application of sentiment analysis in emerging domains such as healthcare, finance, and politics could uncover new challenges and opportunities for research and innovation. Overall, continued research in interpretable sentiment analysis holds the potential to drive meaningful advancements in AI technologies and contribute to more informed decision-making in various fields.
Data Availability
@@
The extended datasets, comprising the Global Twitter COVID-19 Dataset and the USA Twitter COVID-19 Dataset, are publicly available for download from the Extended Covid Twitter Datasets (https://data.mendeley.com/datasets/2ynwykrfgf/1) repository [41]. Additionally, the external datasets used in our research were sourced from Kaggle, including the Twitter and Reddit Dataset (https://www.kaggle.com/datasets/cosmos98/twitter-and-reddit-sentimental-analysis-dataset), Apple Dataset (https://www.kaggle.com/datasets/seriousran/appletwittersentimenttexts), and US Airline Dataset
(https://www.kaggle.com/datasets/crowdflower/twitter-airline-sentiment).